OpenStudy (rockstar0765):

George and carmen went on a bicycle trip. they took a bus to their starting point, and then biked the rest. They traveled 275 kilmoeter in total, and they biked 55 kilometers more than they were bussed. Let x= kilometers traveled by bike and y=kilometers traveled by bus. Find how many kilometer they traveled by bike.

OpenStudy (rockstar0765):

will give medals to whoever answers it right

OpenStudy (quietus):

(275 / 2) + 27.5 = kilometers travelled by bike. (275 / 2) - 27.5 = kilometers travelled by bus.

OpenStudy (rockstar0765):

so what do you mean by this

OpenStudy (quietus):

165km by bike, 110 km by bus.

OpenStudy (anonymous):

Substitute or X and solve for "Y": Y+55 + Y = 275 2Y = 220 Y = 110 km (distance traveled by Bus) ------- Solve for "X": X = Y + 55 X = 110+55 = 165 km (distance traveled by Bike)

OpenStudy (quietus):